# Robust Multi-View Learning via Representation Fusion of Sample-Level Attention and Alignment of Simulated Perturbation

### v1

@@ Abstract
Recently, multi-view learning (MVL) has garnered significant attention due to its ability to fuse discriminative information from multiple views. However, real-world multi-view datasets are often heterogeneous and imperfect, which usually makes MVL methods designed for specific combinations of views lack application potential and limits their effectiveness. To address this issue, we propose a novel robust MVL method (namely RML) with simultaneous representation fusion and alignment. Specifically, we introduce a simple yet effective multi-view transformer fusion network where we transform heterogeneous multi-view data into homogeneous word embeddings, and then integrate multiple views by the sample-level attention mechanism to obtain a fused representation. Furthermore, we propose a simulated perturbation based multi-view contrastive learning framework that dynamically generates the noise and unusable perturbations for simulating imperfect data conditions. The simulated noisy and unusable data obtain two distinct fused representations, and we utilize contrastive learning to align them for learning discriminative and robust representations. Our RML is self-supervised and can also be applied for downstream tasks as a regularization. In experiments, we employ it in unsupervised multi-view clustering, noise-label classification, and as a plug-and-play module for cross-modal hashing retrieval. Extensive comparison experiments and ablation studies validate the effectiveness of RML.
@@ 1 Introduction
In real-world applications, algorithms usually need to handle data with multiple views or modalities in different forms, such as multi-view data from different sensors [74, 36], image-text and video-audio pairs in multimedia [49, 20], and multi-omics features in biomedical data analysis [31]. Compared to a single view, multiple views contain richer information and utilizing them to train more comprehensive machine learning models has given rise to a continuously intriguing research topic, i.e., multi-view learning (MVL). The key to MVL lies in leveraging the explicit correspondences among multiple views for achieving their mutual alignment and information fusion during learning representations, and thus to promote the performance of downstream tasks like clustering [20, 44], classification [52], and retrieval [22].
@@
Despite the significant progress made by previous methods, the following open challenges for MVL still need to be addressed, inspiring us to explore ongoing solutions. Firstly, the heterogeneity of multi-view datasets challenges the universality of MVL methods. Specifically, generalized multi-view data lack a fixed format [77], and there are differences in data modalities, dimensions, sparsity, and scales across heterogeneous views as shown in Figure 1(Left). However, many methods are typically designed with specialized model structures for specific views and modalities [20, 52], making it difficult to apply successful experiences to other applications with different data. For example, the methods designed for multi-view [68] or visual-audio-textual modalities [7] might perform poorly on data with other views as shown in Figure 1(Right). The complexity of real-world applications also makes it nearly impossible to develop specialized models for arbitrary combinations of views. Secondly, real-world multi-view data often are imperfect that contain noise data, unusable data, and noise labels [60, 14], demanding the robustness of MVL methods. Although some methods considered the issue of low-quality multi-view data [79, 63, 58, 76], they primarily focused on balancing a small number of weights at the whole view level rather than addressing it at the finer-grained sample level. Moreover, existing research rarely explores how to design a general MVL method which can enhance the model robustness for multiple different downstream tasks across different learning settings. To address the aforementioned issues, we propose a novel MVL method entitled RML: Robust Multi-View Learning via Representation Fusion of Sample-Level Attention and Alignment of Simulated Perturbation as shown in Figure 2, which enhances the model robustness towards heterogeneous multi-view datasets and possesses the universality by the view-agnostic design to facilitate various downstream tasks.
@@ 2 Method
In this section, we introduce the model framework, training objective, and regularization of RML. We put the related work in Appendix due to space limitations of this paper.
@@ 2.2 Simulated perturbation based multi-view contrastive learning
Concretely, motivated by the success of data augmentation [64, 67, 38], to enhance the model robustness to imperfect multiple views, we randomly add the noise perturbation and the unusable perturbation on partial views to respectively simulate the noisy data and unusable data. The model is trained to achieve the representation alignment before and after dynamic perturbations for eventually resisting them. Formally, we define the simulated perturbations as follows.
@@
where $d({\mathbf{z}}_{i}^{N},{\mathbf{z}}_{i}^{M})=\frac{{\mathbf{z}}_{i}^{N}\cdot{\mathbf{z}}_{i}^{M}}{\|{\mathbf{z}}_{i}^{N}\|_{2}\|{\mathbf{z}}_{i}^{M}\|_{2}}$ measures the distance between two sample representations by cosine similarity. $\tau$ is a controllable temperature parameter in the InfoNCE loss. For $\mathbf{z}_{i}^{N}$, $\mathcal{N}_{i}^{N}=\{{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$ denotes the set of representations to construct the negative sample pairs, i.e., $\{{\mathbf{z}}_{i}^{N},{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$. Similarly, for $\mathbf{z}_{i}^{M}$, $\mathcal{N}_{i}^{M}=\{{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$ and the negative sample pairs are $\{{\mathbf{z}}_{i}^{M},{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$. By minimizing Eq. (13), the multi-view contrastive learning brings similar fused representations closer and pushes dissimilar ones apart. This facilitates the model to capture discriminative information across multi-view data for benefiting downstream tasks.
@@ 3 Main Results
Datasets. Multi-view data is prevalent in real-world applications and exhibits significant heterogeneity. Different datasets usually vary in data modalities, dimensions, sparsity, the number and format of views. Since a MVL method which is compatible with various multi-view datasets is highly anticipated, we conducted experiments on multiple types of multi-view datasets to validate the effectiveness and universality of our method. The information of the used benchmark datasets is shown in Table 2 and the detail is in Appendix [^1] We provide the detail experiment settings and more results of this paper in Appendix..
@@ 3.1 RML on unsupervised multi-view clustering
Settings. In this part, we conduct representation learning and clustering by RML to evaluate its performance of multi-view fusion. To be specific, we first utilize RML to learn fused representations on extensive multi-view datasets (i.e., DHA, BDGP, Prokaryotic, Cora, YoutubeVideo, WebKB, VOC, NGs, and Cifar100), and then perform K-Means on the fused representations to show the clustering quality of RML. The comparison methods include the classical K-Means [39] and eight recent deep learning based multi-view clustering methods: MCN [7], CPSPAN [25], CVCL [8], DSIMVC [59], DSMVC [58], MFLVC [68], SCM [38], and SCMRE [38].
Table 3: Performance comparison on noise-label multi-view classification (with 50% noise label rate and Appendix shows more results)
@@
Table 4: Performance comparison on cross-modal hashing retrieval (16, 32, 64, and 128 represent the different lengths of hash code)
@@ 3.2 RML on noise-label multi-view classification
Settings. In this part, we conduct noise-label multi-view classification to evaluate the robustness of RML against noise labels. Specifically, our experiments are carried out on datasets DHA, BDGP, Prokaryotic, Cora, and YoutubeVideo. We follow the setting of noise-label learning [14] and adopt the symmetric noise labels, which constructs the noise labels for a percentage of training samples by randomly replacing their truth labels with all possible labels. The partition of training set and test set is $7:3$. Our method includes $\rm RML+\mathcal{L}_{\rm CE}$ and $\rm RML+\mathcal{L}_{\rm MCE}$, which optimize the original cross-entropy loss $\mathcal{L}_{\rm CE}$ and multiple cross-entropy losses $\mathcal{L}_{\rm MCE}$ defined in Appendix, respectively. Two baselines $\rm Trans.+\mathcal{L}_{CE}$ and $\rm Trans.+\mathcal{L}_{MCE}$ only minimize $\mathcal{L}_{\rm CE}$ and $\mathcal{L}_{\rm MCE}$, respectively, in which we adopt the same multi-view transformer fusion network as RML for a fair comparison.
Comparison experiments. The performance is evaluated by the commonly-used metrics including classification accuracy (ACC), Precision (Pre.), and F1-score (F1). We report average results of 5 independent runs as shown in Table 3 and have observations as follows: I) From the overall results, we observe that $\rm RML+\mathcal{L}_{\rm MCE}$ and $\rm Trans.+\mathcal{L}_{\rm MCE}$ respectively have better classification results than $\rm RML+\mathcal{L}_{\rm CE}$ and $\rm Trans.+\mathcal{L}_{\rm CE}$ in most cases. Our proposed two simulated perturbations allow us to design new cross-entropy loss $\mathcal{L}_{\rm MCE}$, and optimizing it makes the model more effective to access the category information in imperfect multi-view data. II) Our proposed RML plays a positive role in the model robustness against noise labels. For example, when the noise label rate is $50\%$, $\rm Trans.+\mathcal{L}_{\rm CE}$ has the accuracy of only $0.437$ on BDGP and $\rm RML+\mathcal{L}_{\rm MCE}$ improves it to $0.936$. More results with various noise label rates are shown in Appendix, and we could find that the classification performance of $\rm RML+\mathcal{L}_{\rm MCE}$ and $\rm RML+\mathcal{L}_{\rm CE}$ consistently outperforms that of $\rm Trans.+\mathcal{L}_{\rm MCE}$ and $\rm Trans.+\mathcal{L}_{\rm CE}$. This suggests that our proposed $\mathcal{L}_{\rm RML}$ can be used as a superior regularization term for multi-view classification tasks and it promotes the model robustness towards noise labels.
@@ 4 Ablation Study and Analysis
Table 5: Ablation experiments of noise-label multi-view classification (the noise label rate is set to 50%)
@@
Figure 4: Ablation experiments of cross-modal hashing retrieval on (a) MIRFLICKR-25K and (b) NUS-WIDE.
@@
Figure 3, Table 5, and Figure 4 showcase the ablation study on unsupervised multi-view clustering, noise-label multi-view classification, and cross-modal retrieval tasks, respectively. Concretely, we let $\rm atten.$, $\rm NP$, and $\rm MP$ denote the sample-level attention, noise perturbation, and unusable perturbation, respectively. $\rm base$ denotes the model without these three components and $\rm full$ is the complete RML model. The results indicate that our proposed $\rm atten.$, $\rm NP$, and $\rm MP$ all contributed to improving the $\rm base$ model. This demonstrates that the components in RML are effective and they have the potential to enhance different downstream tasks.
@@
5 Conclusion and Broader Impacts
In the literature, multi-view learning methods have achieved promising progress in fields such as image-text interactions. However, the existing successful experiences are challenging to replicate in the data from like some medical or internet applications, due to the heterogeneous and imperfect nature of multi-view datasets in these areas. In this paper, we propose a novel robust multi-view learning method RML, which is capable of learning fused representations to extract discriminative information from diverse multi-view datasets. Our extensive experiments demonstrate that RML shows promising versatility and it can I) achieve effective multi-view fusion to enhance the unsupervised multi-view clustering, II) increase the model robustness in noise-label multi-view classification, and III) serve as a regularization term to facilitate cross-modal hashing retrieval tasks. Our future work is to extend the framework to more multi-view learning tasks.
                                                                    Appendix
@@ 6.1 Multi-view learning
Multi-view learning (MVL) refers to learning comprehensive information by models from multiple views with matched correspondences. In this paper, we focus on deep learning based MVL methods and categorize existing methods into two types, i.e., representation fusion and representation alignment. Representation fusion methods are the earliest popular in deep MVL, which aims to obtain a fused representation that is superior to representations of individual views [1, 42]. Many of these methods produce more accurate results on the fused representation than that on individual views’ representations, and use it to further refine their representation learning [79, 70]. Representation alignment methods are first investigated by canonical correlation analysis based deep MVL approaches [2, 73, 57]. With the advancement of contrastive learning from self-supervised learning, an increasing number of deep MVL methods have adopted contrastive learning to capture the agreement across views [62, 37, 34, 63, 68]. To achieve the representation alignment, these contrastive MVL methods treat different views of a sample as positive pairs and maximize the similarity among their representations, thereby aiming to learn the semantic information across multiple views [59, 72, 25, 8]. Different from previous deep MVL methods, our RML performs the sample-level attention based multi-view representation fusion, and then achieves the simulated perturbation based representation alignment between the fused representations rather than between views.
@@ 6.2 Contrastive learning
Contrastive learning is a validated and effective paradigm for self-supervised representation learning [13, 51]. It usually constructs positive and negative sample pairs and encourages the model to learn discriminative representations, thereby aggregating the representations of positive sample pairs closer [45, 9]. The approaches for constructing positive sample pairs vary with different types of data. For instance, in terms of image data, data augmentation techniques such as rotation and color filtering are typically employed to generate multiple images that are semantically consistent [17, 12]. For time-series data, adjacent samples in the sequence are used to construct positive sample pairs [46, 48]. Recently, contrastive learning has made significant progress in multi-view or multimodal domains, where different views or modalities of a sample are treated as positive sample pairs without the need for data augmentation [62, 16, 49]. In this work, we propose a novel simulated perturbation based multi-view contrastive learning method for representation learning and downstream tasks, where the positive sample pairs are constructed by the two perturbed versions of fused representations.

### v2

@@ Abstract
Recently, multi-view learning (MVL) has garnered significant attention due to its ability to fuse discriminative information from multiple views. However, real-world multi-view datasets are often heterogeneous and imperfect, which usually causes MVL methods designed for specific combinations of views to lack application potential and limits their effectiveness. To address this issue, we propose a novel robust MVL method (namely RML) with simultaneous representation fusion and alignment. Specifically, we introduce a simple yet effective multi-view transformer fusion network where we transform heterogeneous multi-view data into homogeneous word embeddings, and then integrate multiple views by the sample-level attention mechanism to obtain a fused representation. Furthermore, we propose a simulated perturbation based multi-view contrastive learning framework that dynamically generates the noise and unusable perturbations for simulating imperfect data conditions. The simulated noisy and unusable data obtain two distinct fused representations, and we utilize contrastive learning to align them for learning discriminative and robust representations. Our RML is self-supervised and can also be applied for downstream tasks as a regularization. In experiments, we employ it in multi-view unsupervised clustering, noise-label classification, and as a plug-and-play module for cross-modal hashing retrieval. Extensive comparison experiments and ablation studies validate RML’s effectiveness. Code is available at https://github.com/SubmissionsIn/RML .
footnotetext: Corresponding Authors.
@@ 1 Introduction
In real-world applications, algorithms usually need to handle data with multiple views or modalities in different forms, such as multi-view data from different sensors [78, 37, 49], image-text and video-audio pairs in multimedia [82, 20, 52], and multi-omics features in biomedical data analysis [31, 32]. Compared to a single view, multiple views contain richer information and utilizing them to train more comprehensive machine learning models has given rise to a continuously intriguing research topic, i.e., multi-view learning (MVL). The key to MVL lies in leveraging the explicit correspondences among multiple views for achieving their mutual alignment and information fusion during learning representations, and thus to enhance the performance of downstream tasks like clustering [20, 46], classification [55], and retrieval [22].
@@
Despite the significant progress made by previous methods, the following open challenges for MVL still need to be addressed, inspiring us to explore ongoing solutions. Firstly, the heterogeneity of multi-view datasets challenges the universality of MVL methods. Specifically, generalized multi-view data lack a fixed format [81], and there are differences in data modalities, dimensions, sparsity, and scales across heterogeneous views as shown in Figure 1(Left). However, many methods are typically designed with specialized model structures for specific views and modalities [20, 55], making it difficult to apply successful experiences to other applications with different data. For example, the methods designed for multi-view [72] or visual-audio-textual modalities [7] might perform poorly on datasets comprising other views as shown in Figure 1(Right). The complexity of real-world applications also makes it nearly impossible to develop specialized models for arbitrary combinations of views. Secondly, real-world multi-view data often are imperfect that contain noise data, unusable data, and noise labels [63, 14], which demands the robustness of MVL methods. Although some methods considered the issue of low-quality multi-view data [84, 66, 61, 80], they primarily focused on balancing a small number of weights at the whole view level rather than addressing it at the finer-grained sample level. Moreover, existing research rarely explores how to design a general MVL method which can enhance the model robustness for multiple different downstream tasks across different learning settings. To address the aforementioned issues, we propose a novel MVL method called RML (Robust Multi-View Learning via Representation Fusion of Sample-Level Attention and Alignment of Simulated Perturbation) as shown in Figure 2, which enhances the model robustness towards heterogeneous multi-view datasets and possesses the universality by the view-agnostic design to facilitate various downstream tasks.
@@ 2 Method
In this section, we introduce the model framework, training objective, and regularization function of RML.
@@ 2.2 Simulated perturbation based multi-view contrastive learning
Concretely, motivated by the success of data augmentation [67, 70, 40, 71], to enhance the model robustness to imperfect multiple views, we randomly add the noise perturbation and the unusable perturbation on partial views to respectively simulate the noisy data and unusable data. The model is trained to achieve the representation alignment before and after dynamic perturbations for eventually resisting them. Formally, we define the simulated perturbations as:
@@
where $d({\mathbf{z}}_{i}^{N},{\mathbf{z}}_{i}^{M})=\frac{{\mathbf{z}}_{i}^{N}\cdot{\mathbf{z}}_{i}^{M}}{\|{\mathbf{z}}_{i}^{N}\|_{2}\|{\mathbf{z}}_{i}^{M}\|_{2}}$ measures the cosine similarity between two sample representations. $\tau$ is a controllable temperature parameter in the InfoNCE loss. For $\mathbf{z}_{i}^{N}$, $\mathcal{N}_{i}^{N}=\{{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$ denotes the set of representations to construct the negative sample pairs, i.e., $\{{\mathbf{z}}_{i}^{N},{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$. Similarly, for $\mathbf{z}_{i}^{M}$, $\mathcal{N}_{i}^{M}=\{{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$ and the negative sample pairs are $\{{\mathbf{z}}_{i}^{M},{\mathbf{z}}_{j}^{v}\}_{j\neq i}^{v=N,M}$. By minimizing Eq. (13), the multi-view contrastive learning brings similar fused representations closer and pushes dissimilar ones apart. This facilitates the model to capture discriminative information across multi-view data for benefiting downstream tasks.
@@ 3 Main Results
Datasets. Multi-view data is prevalent in real-world applications and exhibits significant heterogeneity. Different datasets usually vary in data modalities, dimensions, sparsity, the number and format of views. Since a MVL method which is compatible with various multi-view datasets is highly anticipated, we conducted experiments on multiple types of multi-view datasets to validate the effectiveness and universality of our method. The information of the used benchmark datasets is listed in Table 2 showing considerable diversityFor more detailed dataset information, related work, experimental settings, and comparison results, please refer to the Appendix or our Code..
@@ 3.1 RML on unsupervised multi-view clustering
Settings. In this part, we conduct representation learning and clustering by RML to evaluate its performance of multi-view fusion. To be specific, we first utilize RML to learn fused representations on extensive multi-view datasets (i.e., DHA, BDGP, Prokaryotic, Cora, YoutubeVideo, WebKB, VOC, NGs, and Cifar100), and then perform K-Means on the fused representations to show the clustering quality of RML. The comparison methods include the classical K-Means [41] and eight recent deep-learning-based multi-view clustering methods: MCN [7], CPSPAN [25], CVCL [8], DSIMVC [62], DSMVC [61], MFLVC [72], SCM [40], and SCMRE [40].
Table 3: Performance comparison on noise-label multi-view classification with 50% noise label rate (Trans. denotes the transformer network)
@@
Table 4: Performance comparison on cross-modal hashing retrieval (16, 32, 64, and 128 represent the different settings of hash code length)
@@
    In this table, we compare RML+NRCH with NRCH and they are not compared with other cross-modal hashing retrieval methods due to the differences in their experimental settings.
@@ 3.2 RML on noise-label multi-view classification
Settings. In this part, we conduct noise-label multi-view classification to evaluate the robustness of RML against noise labels. Specifically, our experiments are carried out on datasets DHA, BDGP, Prokaryotic, Cora, and YoutubeVideo. We follow the setting of noise-label learning [14, 39] and adopt the symmetric noise labels, which constructs the noise labels for a percentage of training samples by randomly replacing their truth labels with all possible labels. The partition of training set and test set is $7:3$. Our method includes $\rm RML+\mathcal{L}_{\rm CE}$ and $\rm RML+\mathcal{L}_{\rm MCE}$, which optimize the original cross-entropy loss $\mathcal{L}_{\rm CE}$ and multiple cross-entropy losses $\mathcal{L}_{\rm MCE}$ defined in Appendix, respectively. Two baselines $\rm Trans.+\mathcal{L}_{CE}$ and $\rm Trans.+\mathcal{L}_{MCE}$ only minimize $\mathcal{L}_{\rm CE}$ and $\mathcal{L}_{\rm MCE}$, respectively, in which we adopt the same multi-view transformer fusion network as RML for a fair comparison.
Comparison experiments. The performance is evaluated by the commonly-used metrics including classification accuracy (ACC), Precision (Pre.), and F1-score (F1). We report average results of 5 independent runs in Table 3 and have observations as follows: I) From the overall results, we observe that $\rm RML+\mathcal{L}_{\rm MCE}$ and $\rm Trans.+\mathcal{L}_{\rm MCE}$ respectively have better classification results than $\rm RML+\mathcal{L}_{\rm CE}$ and $\rm Trans.+\mathcal{L}_{\rm CE}$ in most cases. Our proposed two simulated perturbations allow us to design new cross-entropy loss $\mathcal{L}_{\rm MCE}$, and optimizing it makes the model more effective to access the category information in imperfect multi-view data. II) Our proposed RML plays a positive role in the model robustness against noise labels. The results show that the classification performance of $\rm RML+\mathcal{L}_{\rm MCE}$ and $\rm RML+\mathcal{L}_{\rm CE}$ consistently outperforms that of $\rm Trans.+\mathcal{L}_{\rm MCE}$ and $\rm Trans.+\mathcal{L}_{\rm CE}$. For example, when the noise label rate is $50\%$, $\rm Trans.+\mathcal{L}_{\rm CE}$ has the accuracy of only $0.437$ on BDGP and $\rm RML+\mathcal{L}_{\rm MCE}$ improves it to $0.936$. This suggests that our proposed $\mathcal{L}_{\rm RML}$ can be used as a superior regularization term for multi-view classification tasks, and the special designs in our framework promote the model’s robustness against noisy labels.
@@ 4 Ablation Study and Analysis
Table 5: Ablation experiments of noise-label multi-view classification with RML+$\mathcal{L}_{\rm MCE}$ (the noise label rate is set to 50%)
@@
Figure 4: Ablation experiments of cross-modal hashing retrieval with RML+UCCH on (a) MIRFLICKR-25K and (b) NUS-WIDE.
@@
Figure 3, Table 5, and Figure 4 present the ablation study on unsupervised multi-view clustering, noise-label multi-view classification, and cross-modal retrieval tasks, respectively. Concretely, we let $\rm atten.$, $\rm NP$, and $\rm MP$ denote the sample-level attention, noise perturbation, and unusable perturbation, respectively. $\rm base$ denotes the model without these three components and $\rm full$ is the complete RML model. The results indicate that our proposed $\rm atten.$, $\rm NP$, and $\rm MP$ all contributed to improving the $\rm base$ model. This demonstrates that the components in RML are effective and they have the potential to enhance different downstream tasks.
@@
5 Conclusion
In the literature, multi-view learning methods have achieved promising progress in fields such as image-text interactions. However, the existing successful experiences are challenging to replicate in the data from like some medical or internet applications, due to the heterogeneous and imperfect nature of multi-view datasets in these areas. In this paper, we propose a novel robust multi-view learning method RML, which is capable of learning fused representations to extract discriminative information from diverse multi-view datasets. Our extensive experiments demonstrate that RML shows promising versatility and it can i) achieve effective multi-view fusion to enhance the unsupervised multi-view clustering, ii) increase the model robustness in noise-label multi-view classification, and iii) serve as a regularization term to facilitate cross-modal hashing retrieval tasks. The future work is to extend the framework to more multi-view learning tasks.
Acknowledgment
This research was supported in part by the National Key Research & Development Program of China under Grant 2022YFA1004100, and in part by the Ministry of Education, Singapore, under its MOE Academic Research Fund Tier 2 (MOE-T2EP20124-0013).
Appendix for Robust Multi-View Learning via Representation Fusion of Sample-Level Attention and Alignment of Simulated Perturbation
@@ 6.1 Multi-view learning
Multi-view learning (MVL) refers to models learning comprehensive information from multiple views with matched correspondences. In this paper, we focus on deep learning based MVL methods and categorize existing methods into two types, i.e., representation fusion and representation alignment. Representation fusion methods are the earliest popular in deep MVL, which aims to obtain a fused representation that is superior to representations of individual views [1, 44]. Many of these methods produce more accurate results on the fused representation than that on individual views’ representations, and use it to further refine their representation learning [84, 74]. Representation alignment methods are first investigated by canonical correlation analysis based deep MVL approaches [2, 77, 60]. With the advancement of contrastive learning from self-supervised learning, an increasing number of deep MVL methods have adopted contrastive learning to capture the agreement across views [65, 38, 35, 66, 72]. To achieve the representation alignment, these contrastive MVL methods treat different views of a sample as positive pairs and maximize the similarity among their representations, thereby aiming to learn the semantic information across multiple views [62, 76, 25, 8]. Different from previous deep MVL methods, our RML performs the sample-level attention based multi-view representation fusion, and then achieves the simulated perturbation based representation alignment between the fused representations rather than between views.
@@ 6.2 Contrastive learning
Contrastive learning is a validated and effective paradigm for self-supervised representation learning [13, 54]. It usually constructs positive and negative sample pairs and encourages the model to learn discriminative representations, thereby aggregating the representations of positive sample pairs closer [47, 9]. The approaches for constructing positive sample pairs vary according to the types of data. For instance, in terms of image data, data augmentation techniques such as rotation and color filtering are typically employed to generate multiple images that are semantically consistent [17, 12]. For time-series data, adjacent samples in the sequence are used to construct positive sample pairs [48, 51]. Recently, contrastive learning has made significant progress in multi-view or multimodal domains, where different views or modalities of a sample are treated as positive sample pairs without the need for data augmentation [65, 16, 52]. Motivated by the success of data augmentation in contrastive learning [67, 70, 40], in this work, we propose a novel simulated perturbation based multi-view contrastive learning method for representation learning and downstream tasks, where the positive sample pairs are constructed by the two perturbed versions of fused representations.